OpenStudy (anonymous):

Find the time required for an investment of 5000 dollars to grow to 7100 dollars at an interest rate of 7.5 percent per year, compounded quarterly. Your answer is t=? :(

OpenStudy (anonymous):

5000=5000*.075*t

OpenStudy (anonymous):

helped??

OpenStudy (anonymous):

well??

OpenStudy (anonymous):

no it wasnt it:( at least for what i solved for..

OpenStudy (dumbcow):

7100 = 5000(1+(.075/4))^4t 7100 = 5000(1.01875)^4t 1.42 = 1.01875^4t ln(1.42)/ln(1.01875) = 4t 18.876 = 4t 4.719 = t

OpenStudy (dumbcow):

you get it?

OpenStudy (anonymous):

actually it said it was wrong..

OpenStudy (anonymous):

your prob close though

OpenStudy (dumbcow):

haha no its right i just verified it t = 4.71912 years

OpenStudy (anonymous):

no it is not right because it said compounded quarterly not continuously. formula is \[5000(1+\frac{.075}{4})^{4y}\]

OpenStudy (anonymous):

then you solve \[7100=5000(1.01875)^{4t}\] \[1.42=(1.01875)^{4t}\] \[\ln(1.42)=4t\ln(1.01875)\] \[t=\frac{\ln(1.42}{4\ln(1.01875)}\]

OpenStudy (anonymous):

oops dumb cow had it i only looked at first one. sorry